OpenStudy (anonymous):

Medal and FAN! How can 1/5x-2 = 1/3x + 8 be set up as a system of equation? 5y − 5x = −10 3y − 3x = 24 5y − 5x = −10 3y + 3x = 24 5y + x = −10 3y + x = 24 5y − x = −10 3y − x = 24

OpenStudy (anonymous):

@chucho78 @jdoe0001 or anyone!

OpenStudy (anonymous):

@aum

OpenStudy (aum):

1/5x-2 = 1/3x + 8 You can set each equal to y and you will have two equations: y = 1/5x - 2 y = 1/3x + 8 multiply the first equation throughout by 5 to get rid of the fraction. multiply the first equation throughout by 3 to get rid of the fraction. In each equation, keep the constant on the right hand side and bring everything else to the left hand side.

OpenStudy (anonymous):

ok thanks!